# Flavor asymmetry from the non-perturbative nucleon sea

### v1

@@ Introduction
Understanding the sea-quark contribution of hadrons is of fundamental importance Kumano:1997cy ; Garvey:2001yq ; Chang:2014jba ; Geesaman:2018ixo . It not only provides a sharper view of the fundamental structure of matter Gottfried:1967kk , but also serves as a tool for detecting physics beyond Standard Model NNPDF:2021njg . Recently, the SeaQuest/E906 experiment published their first measurement of the flavor asymmetry $\bar{d}/\bar{u}$ of the proton sea from the muon pair production within the Drell-Yan process of high-energy proton-ion collisions SeaQuest:2021zxb ; Kotlorz:2021goc ; FNALE906:2022xdu , challenging the previous measurement by the NuSea/E866 experiment at large momentum fraction $x\gtrsim 0.3$ NA51:1994xrz ; HERMES:1998uvc ; NuSea:1998tww ; NuSea:1998kqi ; NuSea:2001idv ; Salajegheh:2017iqp . In particular, the NuSea/E866 experiment predicted that $\bar{d}/\bar{u}$ dips below unity at $x\sim 0.3$, which is contrary to the SeaQuest/E906 measurement. The tension between these experimental measurements calls for a better understanding from theories including global analysis Harland-Lang:2014zoa ; Alekhin:2017kpj ; Accardi:2019ofk ; NNPDF:2021njg ; Cocuzza:2021cbi ; Hou:2022onq ; Cocuzza:2022jye . It was generally believed that the flavor asymmetry of the proton encodes the non-perturbative physics of QCD, making it a particular challenge for calculations from first principles Scapellato:2019csz ; Ji:2020ect .
Among various theoretical interpretations Henley:1990kw ; Dorokhov:1993fc ; Steffens:1996bc ; Londergan:1998ai ; Miller:2002ig ; Trevisan:2008zz ; Song:2011fc ; Kamano:2012yx ; Burkardt:2012hk ; Ji:2013bca ; Ehlers:2014jpa ; Salamu:2014pka ; Lin:2014zya ; Kofler:2017uzq ; Tadepalli:2019ixa ; Bourrely:2019ysc ; Soffer:2019gbb ; Leon:2020bvt ; Alexandrou:2021oih ; Wang:2021ild ; Alberg:2021nmu ; Signal:2021aum ; Cui:2021gzg ; Kotlorz:2021goc ; Chang:2022jri ; Leinweber:2022guz ; He:2022leb ; Bourrely:2022mjf ; Wang:2022bxo ; Bonvini:2023tnk ; Choudhary:2023unw ; Choudhary:2023bap , the pion cloud model Theberge:1980ye ; Thomas:1981vc ; Thomas:1983fh has shown a trend consistent with the SeaQuest/E906 results as well as other theories at small and modest $x$. In this model, the sea quark contribution is viewed as the hadron cloud around the nucleon. The interaction between the pion cloud and the host nucleon is described by chiral effective field theory ($\chi$EFT). Recently, Alberg et. al. showed that light-cone methods can be used to compute the pionic longitudinal momentum distribution function (LMDF) $f_{\pi N}(x)$ Alberg:2012wr ; Alberg:2017ijg ; Alberg:2021nmu . Then, the quark distribution function can be represented as,
@@
To obtain the state vector $|N\rangle_{\text{phy}}$ of the physical nucleon, we diagonalize the light-front quantized Hamiltonian in Fock space. Fig. 1 shows the first few diagrams of the system of equations derived from the light-front Schrödinger equation, which are an infinite set of coupled integral equations in Fock space. A review of the relevant Feynman rules is collected in Refs. Brodsky:1997de ; Carbonell:1998rj . Note that the non-perturbative vertex functions (shaded blocks) are employed to extend the Feynman rules to the non-perturbative regime. The vertex functions are the $T$-matrix elements and are related to the light-front wave functions by resolvents Hiller:1998cv . The explicit expressions translated from these diagrams are shown in Refs. Li:2015iaw ; Karmanov:2016yzu .
@@
While the scalar theory in (3+1)D is super-renormalizable, it still contains divergences. The divergences can be absorbed into the bare parameters by imposing the mass and coupling constant renormalization conditions Li:2015iaw . For the mass renormalization, we require the mass eigenvalue equals the physical mass. For the coupling constant renormalization, we require the $T$-matrix $\Gamma_{N\pi/N}$ equals the physical coupling multiplied by the necessary $Z$ factors at the on-shell point. The whole procedure is non-perturbative. N.B. the values of the bare parameters, i.e., the bare mass $m_{0}^{2}$ and the bare couplings $g_{0}$, depend on the Fock sector truncation. Fortunately, boost invariance in light-front dynamics guarantees that these sector dependent counter-terms can be obtained sector by sector recursively. This renormalization scheme is known as the Fock sector dependent renormalization Karmanov:2008br , which has been successful in renormalizing the light-front Hamiltonian operators of the scalar theory Li:2015iaw ; Karmanov:2016yzu ; Cao:2023ohj , Yukawa theory Karmanov:2010ih ; Karmanov:2012aj and QED Chabysheva:2009vm ; Chabysheva:2009ez as well as $\chi$EFT type interactions Tsirova:2010zza .
@@
To demonstrate the effects of the multi-pion sea, we show in Fig. 3 the LMDFs of the pion inside the physical nucleon for two non-perturbative couplings, $\alpha=1.0$ and $\alpha=2.0$, up to one nucleon and three pions. We also turned off the coupling to $\Delta$, i.e. setting $g_{\Delta}=0$, to focus on the multi-pion sea. Contributions from different Fock sectors are also shown. Comparing these LMDFs, as may be anticipated, the contribution from the multi-pion Fock sector increases with the increase in the coupling.
@@
Figure 3: (Colors online) Longitudinal momentum distribution function of the pion inside the physical scalar nucleon for coupling (Top) $\alpha=1.0$ and (Bottom) $\alpha=2.0$. The coupling to $\Delta$ is turned off in both cases, i.e. $g_{\Delta}=0$. The contributions from the multi-pion Fock sectors are also shown in each panel. For $\alpha=2.0$ there are significant three- and four-body contributions at the large longitudinal momentum fraction $x$.
@@
To compare with the Drell-Yan data, we adopt $g_{N\pi}=6.0\,\mathrm{GeV}$ ($\alpha=0.8$) and $g_{\Delta}=6.0\,\mathrm{GeV}$. These parameters are chosen to reproduce the $\bar{d}-\bar{u}$ data. Fig. 4 shows the LMDFs of the nucleon and $\Delta$ obtained from the two-body, three-body and four-body truncations. As we mentioned, the two-body truncation is equivalent to the light-cone perturbation theory. The results obtained from three- and four-body truncations are close to each other, indicating a trend of numerical convergence of the Fock sector expansion. Thereafter, we will adopt the four-body truncation as the non-perturbative result, and estimate the uncertainty using the difference between the three-body and four-body truncations. Alberg et. al.’s perturbative results based on a chiral Lagrangian are also shown in Fig. 4 for comparison Alberg:2021nmu . In general, our perturbative results differ from their perturbative results since we adopt the scalar version of the theory. Nevertheless, both the perturbative results of Ref. Alberg:2021nmu (not shown in Fig. 4) and our non-perturbative results reproduce the NuSea data $\bar{d}-\bar{u}$ as shown in Fig. 5. The integrated flavor asymmetry $\int{\mathrm{d}}x[\bar{d}(x)-\bar{u}(x)]=0.070(2)$ is consistent with the NuSea/E866 measurement 0.080(11) within the range $0.350.015<x<$ 0.35 NuSea:2001idv .
@@
Figure 4: (Colors online) Longitudinal momentum distribution function of the nucleon (Top) and $\Delta$ (Bottom) inside the physical nucleon. Results from Fock sector truncation up to one baryon plus one pion ($n=2$), one baryon plus two pions ($n=3$) and one baryon plus three pions ($n=4$) are compared. The two-body truncation is equivalent to LO light-cone perturbation theory. Alberg et. al.’s results using LO perturbation theory with a chiral Lagrangian are included for comparison Alberg:2021nmu .
@@
Figure 5 shows the flavor asymmetry of the “scalar nucleon” sea from this theory along with available experimental data. While perturbative and non-perturbative results for $\bar{d}-\bar{u}$ are close to each other, their difference in $\bar{d}/\bar{u}$ is moderately large, which signifies the high Fock sector contributions in the non-perturbative regime especially at large $x$. Within the same theory, our three- (one baryon plus up to two pions) and four-body (one baryon plus up to three pions) truncation results show a good convergence pattern as seen in Fig. 4. However, these results are different from the two-body truncation results. The flavor asymmetry results exhibit significant sensitivity to the 2-body versus 4-body truncation as shown in Fig. 5, which signals important non-perturbative effects originating from the multi-pion sea contributions.
@@
Figure 5: Flavor asymmetry of the nucleon sea: (Top) $\bar{d}-\bar{u}$; (Bottom) $\bar{d}/\bar{u}$ from the perturbative and non-perturbative solutions of a pion cloud model with scalar type interactions in comparison with available experimental data from NuSea/E866 NuSea:1998kqi ; NuSea:2001idv ; Salajegheh:2017iqp and SeaQuest/E906 SeaQuest:2021zxb . The non-perturbative result is obtained with a Fock space truncation up to four-body (one baryon plus up to three pions). The difference between the results with three-body and four-body truncations are shown as bands to indicate the convergence of the Fock space expansion.
@@ Acknowledgements
The authors acknowledge valuable discussions with W. Du, V.A. Karmanov and P. Maris. This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12375081, by the Chinese Academy of Sciences under Grant No. YSBR-101, and by the US Department of Energy (DOE) under Grant No. DE-SC0023692.

### v2

# thanks: Corresponding author

@@ Introduction
Understanding the sea-quark contribution of hadrons is of fundamental importance Kumano:1997cy ; Garvey:2001yq ; Chang:2014jba ; Geesaman:2018ixo . It not only provides a sharper view of the fundamental structure of matter Gottfried:1967kk , but also serves as a tool for detecting physics beyond Standard Model NNPDF:2021njg . Recently, the SeaQuest/E906 experiment published their first measurement of the flavor asymmetry $\bar{d}/\bar{u}$ of the proton sea from the muon pair production within the Drell-Yan process of high-energy proton-ion collisions SeaQuest:2021zxb ; Kotlorz:2021goc ; FNALE906:2022xdu , challenging the previous measurement by the NuSea/E866 experiment at large momentum fraction $x\gtrsim 0.3$ NA51:1994xrz ; HERMES:1998uvc ; NuSea:1998tww ; NuSea:1998kqi ; NuSea:2001idv ; Salajegheh:2017iqp . In particular, the NuSea/E866 experiment predicted that $\bar{d}/\bar{u}$ dips below unity at $x\sim 0.3$, which is contrary to the SeaQuest/E906 measurement. The tension between these experimental measurements calls for a better understanding from theories including global analysis Harland-Lang:2014zoa ; Alekhin:2017kpj ; Accardi:2019ofk ; NNPDF:2021njg ; Cocuzza:2021cbi ; Hou:2022onq ; Cocuzza:2022jye . It was generally believed that the flavor asymmetry of the proton encodes the non-perturbative physics of QCD, making it a particular challenge for calculations from first principles Scapellato:2019csz ; Ji:2020ect . The flavor asymmetry is also related to other sea quark asymmetries, such as the nucleon strangeness as well as the intrinsic charm Brodsky:1996hc .
Among various theoretical interpretations Henley:1990kw ; Dorokhov:1993fc ; Steffens:1996bc ; Londergan:1998ai ; Miller:2002ig ; Trevisan:2008zz ; Song:2011fc ; Kamano:2012yx ; Burkardt:2012hk ; Ji:2013bca ; Wang:2014sua ; Ehlers:2014jpa ; Salamu:2014pka ; Lin:2014zya ; Basso:2015lua ; Kofler:2017uzq ; Tadepalli:2019ixa ; Bourrely:2019ysc ; Soffer:2019gbb ; Leon:2020bvt ; Alexandrou:2021oih ; Wang:2021ild ; Alberg:2021nmu ; Signal:2021aum ; Cui:2021gzg ; Kotlorz:2021goc ; Chang:2022jri ; Leinweber:2022guz ; He:2022leb ; Bourrely:2022mjf ; Wang:2022bxo ; Bonvini:2023tnk ; Choudhary:2023unw ; Choudhary:2023bap , the pion cloud model Theberge:1980ye ; Thomas:1981vc ; Thomas:1983fh has shown a trend consistent with the SeaQuest/E906 results as well as other theories at small and modest $x$. It can also be used to investigate the sea quark-antiquark asymmetry Brodsky:1996hc ; Sufian:2018cpj ; Sufian:2020coz . In this model, the sea quark contribution is viewed as the hadron cloud around the nucleon. The interaction between the pion cloud and the host nucleon is described by chiral effective field theory ($\chi$EFT). Recently, Alberg et. al. showed that light-cone methods can be used to compute the pionic longitudinal momentum distribution function (LMDF) $f_{\pi N}(x)$ Alberg:2012wr ; Alberg:2017ijg ; Alberg:2021nmu . Then, the quark distribution function can be represented as,
@@
To obtain the state vector $|N\rangle_{\text{phy}}$ of the physical nucleon, we diagonalize the light-front quantized Hamiltonian in Fock space. See Refs. Pauli:1985pv ; Pauli:1985ps ; Hornbostel:1988fb ; Tang:1991rc ; Brodsky:1991ir ; Brodsky:1997de ; Vary:2009qz ; Vary:2009gt for reviews of non-perturbative methods developed to diagonalize the light-front Hamiltonian. Fig. 1 shows the first few diagrams of the system of equations derived from the light-front Schrödinger equation, which are an infinite set of coupled integral equations in Fock space. A review of the relevant Feynman rules is collected in Refs. Brodsky:1997de ; Carbonell:1998rj . Note that the non-perturbative vertex functions (shaded blocks) are employed to extend the Feynman rules to the non-perturbative regime. The vertex functions are the $T$-matrix elements and are related to the light-front wave functions by resolvents Hiller:1998cv . The explicit expressions translated from these diagrams are shown in Refs. Li:2015iaw ; Karmanov:2016yzu .
@@
While the scalar theory in (3+1)D is super-renormalizable, it still contains divergences. The divergences can be absorbed into the bare parameters by imposing the mass and coupling constant renormalization conditions Li:2015iaw . For the mass renormalization, we require the mass eigenvalue equals the physical mass. For the coupling constant renormalization, we require the $T$-matrix $\Gamma_{N\pi/N}$ equals the physical coupling multiplied by the necessary $Z$ factors at the on-shell point. The whole procedure is non-perturbative, and is the standard one, to be distinguished with the perturbative treatment in Ref. Alberg:2021nmu . N.B. the values of the bare parameters, i.e., the bare mass $m_{0}^{2}$ and the bare couplings $g_{0}$, depend on the Fock sector truncation. Fortunately, boost invariance in light-front dynamics guarantees that these sector dependent counter-terms can be obtained sector by sector recursively. This renormalization scheme is known as the Fock sector dependent renormalization Karmanov:2008br , which has been successful in renormalizing the light-front Hamiltonian operators of the scalar theory Li:2015iaw ; Karmanov:2016yzu ; Cao:2023ohj , Yukawa theory Karmanov:2010ih ; Karmanov:2012aj and QED Chabysheva:2009vm ; Chabysheva:2009ez as well as $\chi$EFT type interactions Tsirova:2010zza .
@@
To demonstrate the effects of the multi-pion sea, we show in Fig. 3 the LMDFs of the pion inside the physical nucleon for two non-perturbative couplings, $\alpha=0.8$ and $\alpha=2.0$, up to one nucleon and three pions. We also turned off the coupling to $\Delta$, i.e. setting $g_{\Delta}=0$, to focus on the multi-pion sea. Contributions from different Fock sectors are also shown. Comparing these LMDFs, as may be anticipated, the contribution from the multi-pion Fock sector increases with the increase in the coupling.
@@
Figure 3: (Colors online) Longitudinal momentum distribution function of the pion inside the physical scalar nucleon for coupling (Top) $\alpha=0.8$ and (Bottom) $\alpha=2.0$. The coupling to $\Delta$ is turned off in both cases, i.e. $g_{\Delta}=0$. The contributions from the multi-pion Fock sectors are also shown in each panel. For $\alpha=2.0$ there are significant three- and four-body contributions at the large longitudinal momentum fraction $x$.
@@
It is tempting to use nonrelativistic reduction to relate $g_{\pi N}\approx 13$ to the couplings in the scalar theory. However, this procedure produces couplings well beyond the critical coupling for Landau poles Li:2015iaw . To compare with the Drell-Yan data, we adopt $g_{N\pi}=6.0\,\mathrm{GeV}$ ($\alpha=0.8$) and $g_{\Delta}=6.5\,\mathrm{GeV}$. These parameters are chosen to reproduce the SeaQuest $\bar{d}-\bar{u}$ data. Fig. 4 shows the LMDFs of the nucleon, the pion and the $\Delta$ obtained from the two-body, three-body and four-body truncations. As we mentioned, the two-body truncation is equivalent to the light-cone perturbation theory. The results obtained from three- and four-body truncations are close to each other, indicating a trend of numerical convergence of the Fock sector expansion. Thereafter, we will adopt the four-body truncation as the non-perturbative result, and estimate the uncertainty using the difference between the three-body and four-body truncations. Alberg et. al.’s perturbative results based on a chiral Lagrangian are also shown in Fig. 4 for comparison Alberg:2021nmu . In general, our perturbative results differ from their perturbative results since we adopt the scalar version of the theory. Nevertheless, both the perturbative results of Ref. Alberg:2021nmu and our non-perturbative results reproduce the SeaQuest data $\bar{d}-\bar{u}$ as shown in Fig. 5. The integrated flavor asymmetry $\int_{0.13}^{0.45}{\mathrm{d}}x[\bar{d}(x)-\bar{u}(x)]=0.0122(7)$ is consistent with the SeaQuest/E906 measurement 0.0159(60) within the range $0.450.13<x<$ 0.45 FNALE906:2022xdu .
@@
Figure 4: (Colors online) Longitudinal momentum distribution function of the nucleon (Top), the pion (Middle) and the $\Delta$ (Bottom) inside the physical nucleon. Results from Fock sector truncation up to one baryon plus one pion ($n=2$), one baryon plus two pions ($n=3$) and one baryon plus three pions ($n=4$) are compared. The two-body truncation is equivalent to LO light-cone perturbation theory. Alberg et. al.’s results using LO perturbation theory with a chiral Lagrangian are included for comparison Alberg:2021nmu .
@@
Figure 5 shows the flavor asymmetry of the “scalar nucleon” sea from this theory along with available experimental data at the scale $Q^{2}=25.5\,\mathrm{GeV}^{2}$. Results from the pion cloud model by Alberg et. al. Alberg:2021nmu and from the statistical model by Basso et. al. Basso:2015lua are also shown for comparison. While perturbative and non-perturbative results for $\bar{d}-\bar{u}$ are close to each other, their difference in $\bar{d}/\bar{u}$ is moderately large, which signifies the high Fock sector contributions in the non-perturbative regime especially at large $x$. Within the same theory, our three- (one baryon plus up to two pions) and four-body (one baryon plus up to three pions) truncation results show a good convergence pattern as seen in Fig. 4. However, these results are different from the two-body truncation results. The flavor asymmetry results exhibit significant sensitivity to the 2-body versus 4-body truncation as shown in Fig. 5, which signals important non-perturbative effects originating from the multi-pion sea contributions.
@@
Figure 5: Flavor asymmetry of the nucleon sea: (Top) $\bar{d}-\bar{u}$; (Bottom) $\bar{d}/\bar{u}$ from the perturbative and non-perturbative solutions of a pion cloud model with scalar type interactions in comparison with available experimental data from NuSea/E866 NuSea:1998kqi ; NuSea:2001idv ; Salajegheh:2017iqp and SeaQuest/E906 SeaQuest:2021zxb ; FNALE906:2022xdu at the scale $Q^{2}=25.5\,\mathrm{GeV}^{2}$. The non-perturbative result is obtained with a Fock space truncation up to four-body (one baryon plus up to three pions). The difference between the results with three-body and four-body truncations are shown as bands to indicate the convergence of the Fock space expansion. Results from the pion cloud model with light-cone perturbation theory by Alberg et. al. Alberg:2021nmu and from the statistical model by Basso et. al. Basso:2015lua are also shown for comparison.
Finally, as a consistency check, Fig. 6 shows the quark-antiquark asymmetry $d(x)$ vs $\bar{d}(x)$ obtained from this model using the non-perturbative solutions. The CT18 NNLO results are shown for comparison Hou:2019efy .
Figure 6: The quark-antiquark asymmetry $d(x)$ vs $\bar{d}(x)$ obtained from this model using the non-perturbative solutions ($n=3,4$) of the scalar Yukawa model. The $n=3$ and $n=4$ results overlap with each other. The CT18 NNLO results are also shown for comparison Hou:2019efy .
@@ Acknowledgements
The authors acknowledge valuable discussions with W. Du, V.A. Karmanov and P. Maris. We thank the referee for the discuss on the low-$x$ region. This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12375081, by the Chinese Academy of Sciences under Grant No. YSBR-101, and by the US Department of Energy (DOE) under Grant No. DE-SC0023692. X. Z. is supported by new faculty startup funding by the Institute of Modern Physics, Chinese Academy of Sciences, by Key Research Program of Frontier Sciences, Chinese Academy of Sciences, Grant No. ZDBS-LY-7020, by the Natural Science Foundation of Gansu Province, China, Grant No. 20JR10RA067, by the Foundation for Key Talents of Gansu Province, by the Central Funds Guiding the Local Science and Technology Development of Gansu Province, Grant No. 22ZY1QA006, by Gansu International Collaboration and Talents Recruitment Base of Particle Physics (2023-2027), by International Partnership Program of the Chinese Academy of Sciences, Grant No. 016GJHZ2022103FN, by National Natural Science Foundation of China, Grant No. 12375143, by National Key R&D Program of China, Grant No. 2023YFA1606903 and by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB34000000.